OpenStudy (anonymous):

Prove that d/dc arcsecu=1/u sqrt u^2-1 du/dx

terenzreignz (terenzreignz):

Two things, the first, and more important being... what's the derivative of the arcsecant function?

terenzreignz (terenzreignz):

\[\huge \frac{d}{dx}\sec^{-1}x\]

OpenStudy (anonymous):

yup.

terenzreignz (terenzreignz):

To understand this, let's look at this function... \[\huge y = \sec^{-1}x\]

OpenStudy (anonymous):

okie

terenzreignz (terenzreignz):

Now, this is (pretty much) the same as \[\huge \sec \ y = x\] Agreed?

OpenStudy (anonymous):

i agree, :)

terenzreignz (terenzreignz):

Well, then, use implicit differentiation, and find dy/dx.

terenzreignz (terenzreignz):

Are you still with me?

OpenStudy (anonymous):

yes lol just trying to remember how to do implicits.

terenzreignz (terenzreignz):

Okay. Keep me posted :)

OpenStudy (anonymous):

how do i start it again? lol

terenzreignz (terenzreignz):

You differentiate x as normal. You differentiate y as you would with x, but when you're done, multiply it by dy/dx

OpenStudy (anonymous):

im confused lol.

OpenStudy (anonymous):

x would just be 1 then?

terenzreignz (terenzreignz):

That's right. As for y, you treat it as a function of x.

terenzreignz (terenzreignz):

And you apply the chain rule. Differentiate y as if it were an x, and afterwards, when the dust settles, multiply the differentiated expression by dy/dx, or y'

OpenStudy (anonymous):

so i got sec y * y' = 1?

terenzreignz (terenzreignz):

sec y? I said differentiate it first, before you multiply the y' :P

OpenStudy (anonymous):

so secytany * y'?

terenzreignz (terenzreignz):

Bingo :) Let me write that down for you... After differentiating both sides with respect to x... we get \[\huge (\sec \ y \ \tan \ y)\frac{dy}{dx}=1\] Ok? :P

OpenStudy (anonymous):

yup!

terenzreignz (terenzreignz):

Now, solve for dy/dx. What I mean is that make it so that dy/dx is alone on one side of the equation.

OpenStudy (anonymous):

dy/dx=1/secytany

terenzreignz (terenzreignz):

Very good :) \[\huge \frac{dy}{dx}=\frac{1}{\sec \ y \ \tan \ y }\]

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Now, earlier, we started with \[\huge y = \sec^{-1}x\] right?

OpenStudy (anonymous):

oui

terenzreignz (terenzreignz):

Well, that's the same as \[\huge \sec \ y = x\] yes?

OpenStudy (anonymous):

yuppp

terenzreignz (terenzreignz):

Well, then, put that to one side, and let's get to work on this... \[\large \frac{dy}{dx}=\frac1{\sec \ y \ \tan \ y}\] Remember the pythagorean identities? Specifically the one that relates secant to tangent?

OpenStudy (anonymous):

sec=h/o?

OpenStudy (anonymous):

t=o/a

terenzreignz (terenzreignz):

No... this identity... \[\large \cos^2\theta+\sin^2\theta=1\] Remember it?

OpenStudy (anonymous):

oya

terenzreignz (terenzreignz):

Well, related to it is the identity involving the tangent and the secant function. Remember it?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

What is it?

OpenStudy (anonymous):

sec^2=1+tan^2

terenzreignz (terenzreignz):

Okay, very good :) \[\large \sec^2x=1+\tan^2x\] So it follows that \[\huge \sec^2x-1=\tan^2x\] catch me so far?

OpenStudy (anonymous):

yup.

terenzreignz (terenzreignz):

Ok, we take the square root of both sides, we get... \[\huge \sqrt{\sec^2x-1}=\tan \ x\] Got it?

OpenStudy (anonymous):

yupp!

terenzreignz (terenzreignz):

So, back to this... \[\huge \frac{dy}{dx}=\frac1{\sec \ y \ \tan \ y}\] Due to the conclusion we just arrived at... We can replace tan y with... \[\huge \frac{dy}{dx}=\frac1{(\sec \ y)\sqrt{\sec^2y-1}}\] Getting it?

OpenStudy (anonymous):

now i just sub back in x because secy=x

terenzreignz (terenzreignz):

You catch on really quickly. So you're finally left with...?

OpenStudy (anonymous):

dy/dx=1/xsqrtx^2-1

terenzreignz (terenzreignz):

Et voila... Well done. With this, you should have no trouble at all answering your question, just one application of chain rule, and you're all set :D

OpenStudy (anonymous):

so to apply the fchain rule, i should just write the dy/dx at the end?

terenzreignz (terenzreignz):

Not dy/dx but du/dx

OpenStudy (anonymous):

yeah thats where im a little confused, how to make it du/dx because thaqts what the question asks for.

terenzreignz (terenzreignz):

Well, could you restate your question? I suspect a typo in your post...

OpenStudy (anonymous):

there is also a second part to this question: use the formula we just came up with in a to find the derivative of y=arcsec(x/a) where a is a constant.

OpenStudy (anonymous):

d/dx arcsecu=1/u squrt u^2-1 du/dx

terenzreignz (terenzreignz):

Chain rule. You just get the derivative of the arcsecant, except replace x with x/a And then multiply it to the derivative of x/a, which is just...?

OpenStudy (anonymous):

i have to run to class! ill be back in about an hr or so. thanks for the help and explanation, it really helped me a lot. : )

OpenStudy (anonymous):

back!